Find parametric equations for the rectangular equation x² + y² = 4.
x = cos(2t), y = sin(2t), 0° < t < 360°
x = 4 cos(t), y = 4 sin(t), 0° < t < 360°
x = 2 cos(t), y = 2 sin(t), 0° < t < 360°
x = 16 cos(t), y = 16 sin(t), 0° < t < 360°

Answer by ikleyn(52781)

.
Find parametric equations for the rectangular equation x² + y² = 4.

x = cos(2t), y = sin(2t), 0° < t < 360°
x = 4 cos(t), y = 4 sin(t), 0° < t < 360°
x = 2 cos(t), y = 2 sin(t), 0° < t < 360°       <--- This and only this
x = 16 cos(t), y = 16 sin(t), 0° < t < 360°